OpenStudy (anonymous):

Evaluate the expression for the given value of the variable. -(4h-1)/1+h ; h = 2 I know I have to plug in the value of h but how do I solve the rest of the equation?

OpenStudy (aum):

It is not an equation and so there is nothing to solve. You just need to evaluate the expression -(4h-1)/1+h when h = 2. Put h = 2 and you get: -(4 * 2 - 1) / (1 + 2) = -7 / 3. I am assuming the expression is: \(\Large -\frac{4h-1}{1+h}\)

OpenStudy (anonymous):

Yes you are correct, that is what it looks like. I thought I had to solve it. Thank you for helping!

OpenStudy (aum):

You are welcome.

OpenStudy (ray1998):

I got up until this point too : ) so thank you for explaining it. My question is, do you carry out by simplifying -7/3 or do you just leave it as it is?